i need to know how to solve for X in this formula, -4(3-x)=2(x+6)
Answer by rapaljer(4671)

-4(3-x)=2(x+6)

Step 1: Remove parentheses by Distributive Property
-12+4x = 2x + 12

Step 2: Get all the x terms on one side by subtracting 2x from each side
-12 + 4x - 2x = 2x -2x + 12
-12 + 2x = 12

Step 3: Get all number terms on the other side by adding +12 to each side
-12 + 12 + 2x = 12 + 12
2x = 24

Step 4: Divide both sides by 2
x=12


Please see my website for more explanations by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Look for "Basic, Intermediate, and College Algebra: One Step at a Time", and choose "Basic Algebra", then "Chapter 1" and Equation Solving or other topic you need help with. See also, my "Math in Living Color" pages.

R^2

Dr. Robert J. Rapalje
Seminole Community College
Altamonte Springs, FL